CALCULATE THE MOLE FRACTION OF EACH COMPONENT OF A SOLUTION CONTAINING 65 G OF ETHANOL IN 350 G OF WATER.


1
Expert's answer
2022-03-27T10:31:31-0400

χi = ni / Σni

ni - moles of component

Σni - total mole of solution


n(C2H5OH) = m(C2H5OH)/Mr(C2H5OH)

n(C2H5OH) = 65 g / 46 g/mole= 1.41 mole


n(H2O) = m(H2O)/Mr(H2O)

n(H2O) = 350 g / 18g/mole = 19.44 mole


Σni = n(C2H5OH) + n(H2O)

Σni = 1.41 mole + 19.44 mole = 20.85 mole


χ(C2H5OH) = n(C2H5OH) / Σni = 1.41 mole / 20.85 mole = 0.068

χ(H2O) = n(H2O) / Σni = 19.44 mole / 20.85 mole = 0.932